### v1

# Monocular Facial Appearance Capture in the Wild

Abstract
We present a new method for reconstructing the appearance properties of human faces from a lightweight capture procedure in an unconstrained environment. Our method recovers the surface geometry, diffuse albedo, specular intensity and specular roughness from a monocular video containing a simple head rotation in-the-wild. Notably, we make no simplifying assumptions on the environment lighting, and we explicitly take visibility and occlusions into account. As a result, our method can produce facial appearance maps that approach the fidelity of studio-based multi-view captures, but with a far easier and cheaper procedure.
1 Introduction
3D facial scanning is a fundamental tool for the creation of realistic digital humans in several industries like film and video game entertainment, communication and telepresence, medical applications, and the new trend of AI-driven digital characters. For decades, practitioners have relied on high-quality 3D face scans in order to bring people into virtual worlds. Much of the technology evolution has focused on reconstructing the surface geometry, where initial scanners could create detailed triangle meshes in controlled studio settings with many cameras and lights, and then more recent efforts focused on fast and lightweight face reconstruction from monocular inputs in less-constrained, so-called “in-the-wild” settings. While the facial geometry is extremely important, the shape alone is not enough to re-render the subject in novel environments with photorealistic quality. For this task, we must additionally recover the appearance properties of the face, which dictate how light interacts with the skin surface. As such, in today’s high-end facial scanning pipelines the desired result includes a high-resolution facial surface mesh with corresponding appearance textures for properties like the diffuse albedo, specular intensity and specular roughness, which are compatible with modern skin shaders.
Like facial geometry reconstruction, the field of skin appearance estimation is also well-studied in controlled studio environments, where accurate appearance maps can be reconstructed from large setups that obtain multi-view or multi-shot images under calibrated lighting. Following the geometry trend, current research aims to allow facial appearance capture in less constrained settings, for example outdoors using the sun as a single point light [57]. Unfortunately, these methods often make simplifying assumptions, and thus there still exists a large gap in reconstruction quality between current in-the-wild methods and production-ready studio-based capture.
In this work we present a new method for facial appearance capture in the wild, surpassing the level of fidelity of existing lightweight methods. Our approach requires only a short video sequence of a simple head rotation, captured from a single camera in any environment, including indoors or outdoors, on a sunny day or in shadow. Our approach is built on traditional inverse rendering optimization, where a fast differentiable renderer is used to solve for the geometry and appearance parameters together with the environment lighting simultaneously. Different from previous methods, we do not make any assumptions on the lighting condition (e.g. we do not require a sun in the sky), and we explicitly take visibility into account, effectively removing baked-in shading by correctly modeling self-occlusion in our appearance solver. We do so by proposing a novel shading model that combines ray tracing and a visibility modulated pre-filtered environment map. As we will show, our approach leads to more faithful recovery of the appearance properties than existing techniques in the wild. The result is a detailed geometry mesh with textures for diffuse albedo, specular intensity and roughness, which approach the quality achievable from studio-based setups but are captured more flexibly in the wild. As a particular application, our approach allows fast capture of actors on a film set, with resulting assets that can be used directly in traditional VFX pipelines. In summary, we make the following key contributions:
  • A new state-of-the-art method for in-the-wild facial appearance capture that makes no assumption on the scene lighting condition.
  • A novel shading model that explicitly handles self-occlusion by combining ray tracing with pre-filtered environment maps, achieving high-quality appearance reconstruction from monocular input.
2 Related Work
In the following section, we first outline relevant works around in-the-wild inverse rendering which do not necessarily focus on the human face. Second, we highlight works which specifically tackle facial appearance capture.
Inverse rendering in the wild.
Inverse rendering is the process of decomposing the scene into 3D shape, material and illumination by simulating the rendering process and comparing the results against captured images. It has been a popular research topic with the recent advances in novel view synthesis using neural implicit representations [44, 45, 56], 3D Gaussian splatting [35], mesh-based differentiable renderers (rasterizers [36] and path tracers [38, 29]) and using diffusion models as prior [41, 21, 39, 60]. While some existing techniques target a very challenging scenario where the lighting can differ across different images [9, 10, 20], we restrict our discussion here to methods that assume a static unknown environment lighting. PhySG [65] utilizes spherical Gaussians to approximately and efficiently evaluate the rendering equation in closed form. Munkberg et al. [46] propose an efficient end-to-end framework for joint learning of topology, triangle meshes and materials, achieving much faster training and inference compared to previous NeRF-based factorization methods [67, 8, 55]. They also introduce a differentiable formulation of the split-sum approximation of environment lighting to efficiently recover all-frequency lighting. Follow-up work [28] shows that material and lighting decomposition can be further improved with a more realistic shading model, incorporating ray tracing and Monte Carlo integration. Recently, 3D Gaussian Splatting techniques were used in conjunction with physical-based rendering to allow for scene relighting [40, 23, 6, 69, 59].
Facial appearance capture.
Traditional face capture studios often employ a multi-view setup with controlled and calibrated lighting conditions to reconstruct the skin appearance [52, 25, 26, 15]. Similar setups were used for facial appearance decomposition with neural and gaussian primitives [53, 64, 62, 61, 1]. Lighter alternative setups were also explored by Lattas et al. [37] and Choi et al. [14] but they still require the subject to be seating in a dedicated space. Recently, the research community has investigated more lightweight setups which are easily accessible to everyone. However, most of these techniques still pose some constraints on the capture environment. CoRA [27] reconstructs relightable 3D face assets from a single co-located smartphone flashlight sequence captured in a dim room. Similarly, Azinović et al. [2] additionally attach polarization foils and capture two such sequences with perpendicular polarization orientation to separate skin surface and subsurface reflectance. Using a co-located light and camera setup, these methods assume the position of the dominant light source is known and no shadowing term needs to be modeled. Instead of using a smartphone flashlight, SunStage [57] takes a selfie video rotating under the sun as input and uses the varying angles between the sun and the face as guidance. Cast shadows from the sun are modeled by shadow mapping. It also jointly optimizes the sun’s position together with face geometry and appearance. All these methods assume the specular reflection (and the shadowing) of the face comes from a single dominant point or directional light source in the capture environment, and the ambient light contributes only to a low-frequency diffuse term. NextFace [16, 17, 18] exploits a differentiable ray tracer to estimate various scene attributes given unconstrained monocular images. It relies heavily on statistical shape and appearance priors [7, 24, 54] which limits its expressiveness. Our work instead focuses on accurate personalized reconstruction. Rainer et al. [51] use tiny shading networks to disentangle shading from explicit reflectance maps but assume a known high-quality geometry and smooth lighting. NeuFace [68] represents the face with a neural SDF and proposes to learn appearance factorization under unknown low frequency light with a novel neural BRDF basis. However, it requires multi-view input similar to those from a light stage setup. Closest to ours, FLARE [5] builds relightable head avatars from monocular videos. It uses pre-filtered environment maps for relighting and a neural approximation of it during training. This rendering model ignores the self-occlusion and bakes part of the shading into the albedos. In contrast, we explicitly handle light visibility, using ray tracing for diffuse reflectance and visibility modulated pre-filtered environment map for the specular reflectance, leading to higher quality shape and appearance reconstruction.
3 Monocular Appearance Capture
We now describe our method for facial appearance capture given a monocular head rotation sequence. We assume the expression of the subject does not change throughout the sequence. As a pre-processing step, we run monocular tracking based on landmarks [12] and a photometric loss [50] to obtain an initial canonical mesh from a 3DMM fit. We also estimate a fixed camera pose, per-frame rigid head poses and neck rotation. For more details about the pre-processing step please see the supplemental PDF. The output of our inverse rendering system will be a 3D asset of the subject containing a high-quality mesh, and diffuse albedo, specular intensity and roughness as 2D texture maps. In the following, we first describe our geometry optimization formulation in Section 3.1, shading model in Section 3.2, and then implementation details in Section 3.3.
3.1 Geometry Optimization
Balancing the updates to both geometry and textures in inverse rendering can be a challenging task. More specifically, the optimization might overfit too quickly on the textural components before learning the correct geometry. This is often the case when Laplacian shape regularization is applied to enforce geometric smoothness, making the geometry update too slow. Related work such as FLARE [5] addresses this issue by using a two-stage approach. First, a detailed geometry and only blurry textures are learned, and then, the shape is fixed and textures are learned after re-initialization in the second stage. Our method, however, optimizes geometry and textures at the same time. To do so, we adopt a preconditioning framework similar to Nicolet et al. [48], which biases gradient steps towards smooth solutions. The vertex positions $\mathbf{v}$ in each iteration are updated by
$\mathbf{v}\leftarrow\mathbf{v}-\eta(\mathbf{I}+\lambda_{\text{geo}}\mathbf{L})^{2}\frac{\partial\mathcal{L}}{\partial\mathbf{v}},$ (1)
where $\mathbf{v}\in R^{N\times 3}$ collects mesh vertex positions along rows, $\eta$ is the learning rate, $\mathbf{I}$ is an identity matrix, $\mathbf{L}$ is the uniform Laplacian, and $\mathcal{L}$ is our loss function described in Section 3.3. The hyper-parameter $\lambda_{geo}>0$ balances between the original objective of matching the input images and a smooth mesh. We set $\lambda_{geo}=19$ in all our experiments. This way, we can apply a large learning rate for the geometry optimization while keeping the mesh smooth and self-intersection free.
3.2 Shading Model
Following the rendering equation [30], we compute the outgoing radiance $L(\omega_{o})$ in direction $\omega_{o}$ by:
[…]
We decompose the BRDF $f(\mathbf{x},\omega_{i},\omega_{o})$ into a diffuse term $f_{d}$ and a specular term $f_{s}$. We use the simple Lambertian model for the diffuse term:
$f_{d}(\mathbf{x})=\frac{\rho(\mathbf{x})}{\pi},$ (3)
where $\rho$ is the diffuse albedo. We use a specular BRDF similar to Kelemen and Szirmay-Kalos [34], which has been shown to be well suited for rendering human skin [47, Chapter 14]:
$f_{s}(\mathbf{x},\omega_{i},\omega_{o})=\frac{DGF}{4(\omega_{i}\cdot\mathbf{n})(\omega_{o}\cdot\mathbf{n})},$ (4)
where D, G, and F are functions representing the Beckmann normal distribution, geometric attenuation and Fresnel terms, respectively.
Accounting for light visibility.
We consider direct illumination where $L_{i}(\omega_{i})$ comes only from the light sources. Spherical Harmonics [42, 13] or Spherical Gaussians [63] are popular representations used by prior methods [65, 68] but they can only model low- to medium-frequency lighting. Other work [5, 46] uses pre-filtered environment maps [32, 43], which allows to capture all-frequency lighting but still ignores self-shadowing. We aim to use a more realistic rendering model that accounts for light visibility to achieve higher photorealism. For the diffuse term, we apply explicit ray tracing to compute the outgoing radiance and use
[…]
where $V(\mathbf{x},\omega_{i})$ is the visibility at surface point $\mathbf{x}$ from incoming light direction $\omega_{i}$. We perform Monte Carlo integration with multiple importance sampling similar to Hasselgren et al. [28] skipping the denoising step as we did not observe any noise issue for the diffuse term in our experiments.
We empirically found that ray tracing for the specular term tends to produce flickery renders (Figure 6) due to our challenging scenario where only one camera view is available. Therefore, we propose to use a pre-filtered environment map but modulate the specular render with a view-dependent visibility term. We apply a split-sum approximation similar to Karis [31] and use
$L_{\text{specular}}(\omega_{o})\approx$ (6)
[…]
The first term is the material integral that depends on $\omega_{i}\cdot\mathbf{n}$ and the roughness. The second integral is the pre-filtered environment map. Both integrals can be pre-computed and stored as 2D textures. At rendering time, the rendering integral can be evaluated using only two texture lookups. When light visibility is not accounted for, the second integral only depends on the reflected view direction $\omega_{r}$ and the roughness $r$. When $r\ll 1$, the second integral is 0 unless $\omega_{i}$ is close to $\omega_{r}$. Therefore, we can approximate it as
Figure 1: A selection of facial appearance reconstruction and decomposition results for different subjects in different environments, both indoors and outdoors, with sunny and cloudy sky.
[…]
Here, $\tilde{V}(\mathbf{x},\omega_{r})$ is the view-dependent specular visibility. The simplest choice of $\tilde{V}(\mathbf{x},\omega_{r})$ is to only evaluate the light visibility at $\omega_{r}$, i.e., $\tilde{V}(\mathbf{x},\omega_{r})\coloneqq V(\mathbf{x},\omega_{r})$. Intuitively, this means the specular component is 0 if the reflected view direction is occluded. Note that this approximation is exact for perfect specular reflection (i.e. mirrors). To add some softness in the visibility attenuation term (instead of considering it as a binary function), we approximate it by Monte Carlo integration,
$\tilde{V}(\mathbf{x},\omega_{r})\coloneqq\frac{1}{K}\sum_{k=1}^{K}\frac{V(\mathbf{x},\omega_{k})}{p(\omega_{k})},$ (8)
where the samples are drawn following the normal distribution of the specular lobe. This bears some resemblance to the appearance models of Saito et al. [53]. However, they use Spherical Gaussians for the specular lobe and parameterize the view-dependent specular visibility term using a neural network as they work with a large amount of studio data where the lighting is controlled and calibrated.
3.3 Implementation Details
Our training loss consists of an $L_{1}$ data term $\mathcal{L}_{\text{img}}$ and some regularization terms. We employ an $L_{1}$ mask loss $\mathcal{L}_{\text{mask}}$ between the mask obtained from MODNet [33] and the predicted binary mask. Although we use a parameterization similar to Nicolet et al. [48], we find it helpful to still employ a Laplacian regularizer to stabilize the geometry optimization such that we do not need to set different $\lambda_{\text{geo}}$ for different datasets. This regularizer encourages the Laplacian of the optimized mesh to stay close to the Laplacian of the initial 3DMM fit
$\mathcal{L}_{\text{Lap}}=\big{\|}\mathbf{L}(\mathbf{v}-\mathbf{v}_{\text{init}})\big{\|}^{2}_{2}.$ (9)
We apply a white light regularization $\mathcal{L}_{\text{light}}$ on the environment map as in [46] and the roughness texture is regularized to be smooth via a total variation loss $\mathcal{L}_{\text{rough}}$. We noticed in our experiments that part of the specular signal tends to be baked into the diffuse albedo. We thus apply a weak regularization to encourage the diffuse render $I_{\text{diffuse}}$ to be small if possible, as
$\mathcal{L}_{\text{diffuse}}=\big{\|}I_{\text{diffuse}}\big{\|}_{2}^{2}.$ (10)
The final loss is then
$\mathcal{L}\coloneqq$ $\mathcal{L}_{\text{img}}+\lambda_{\text{mask}}\mathcal{L}_{\text{mask}}+\lambda_{\text{Lap}}\mathcal{L}_{\text{Lap}}+$ (11)
$\lambda_{\text{light}}\mathcal{L}_{\text{light}}+\lambda_{\text{rough}}\mathcal{L}_{\text{rough}}+\lambda_{\text{diffuse}}\mathcal{L}_{\text{diffuse}},$
where the weights are defined as
$\lambda_{\text{mask}}=0.1,\lambda_{\text{Lap}}=10,\lambda_{\text{light}}=0.1,$ (12)
$\lambda_{\text{rough}}=0.1,\lambda_{\text{diffuse}}=0.01.$
We capture 500 to 800 frames for each subject. Please see the supplemental PDF for details of our capture protocol. We then uniformly sample around 250 frames to use in the inverse rendering. We did not observe a performance gain or drop when using all of the frames. The images are cropped to 1K resolution. We also solve the texture maps at 1K resolution. The environment map is a cubemap with resolution $6\times 256\times 256$ at the largest mip-level, and resolution $6\times 8\times 8$ at the smallest mip-level. For each pixel, we draw $256$ light samples and $256$ BRDF (cosine) samples for the diffuse render, and $64$ samples to estimate the view-dependent specular visibility. We employ the Adam optimizer [19] with a learning of $0.1$ for vertex positions and the environment map, and $0.001$ for the textures. We use the differentiable rasterizer from Laine et al. [36] to obtain the primary visibility and the OptiX [49] engine for ray tracing. Each subject is trained for $6000$ iterations which takes around 2 hours on a Nvidia RTX 3090 GPU.
4 Experiments
We now show several results of our appearance capture method, evaluate its performance compared to previous work, and offer several ablation studies to validate our design choices. Please refer to our supplemental material for additional results.
4.1 Appearance and Geometry Reconstruction
We begin by highlighting the versatility of our approach by showing several appearance capture results of different subjects in different environments in Fig. 1. Each row of the figure illustrates one of the input images, the corresponding render using our recovered appearance properties, and then a breakdown of the reconstructed appearance maps (diffuse albedo, specular intensity, specular roughness) and geometry. Our method can be applied indoors or outdoors, with sunny or cloudy skies. We show how the recovered geometry and appearance maps can be used to render the subject in a new environment by relighting them. Fig. 2 shows two subjects relit in two different environments (see Fig. 1 for the original environment).
Figure 2: Relighting results of multiple frames on two different subjects in two different environments. The environment is shown top left of each row. See Fig. 1 for the original environment.
4.2 Comparisons
We compare our results to related methods for facial appearance capture in the wild: FLARE [5], NextFace [16], and SunStage [57].
Fig. 3 shows a qualitative comparison of our method to FLARE on two different subjects. While the combined final render is similar between FLARE and our method, it is clear that FLARE fails to separate the diffuse and specular components, baking most of the specular signal in the diffuse map resulting in a nearly zero specular render (Fig. 3, 3rd column). In contrast, our shading model is completely physically-based and correctly separates the diffuse and specular components. The normals reconstructed by FLARE portray a lot of spatial noise and the geometry contains self-intersections, unlike ours (4th column). The diffuse albedo from FLARE contains more baked-in diffuse shading than our result (5th column). As a result, FLARE performs worse when relighting the subject under a novel environment map than our approach (columns 6, 7 and 8).
Figure 3: Qualitative comparisons with FLARE [5]. We show one training frame on the left, with separated diffuse, specular, normals and diffuse albedo renders. On the right we show a relit example using the reconstructed appearance. FLARE fails to separate the diffuse and specular signals, produces noisy normals, and bakes shading into the diffuse albedo, all leading to poor results under relighting.
We also perform qualitative comparisons to NextFace and SunStage in Fig. 4. NextFace relies on a statistical prior, leading to inaccurate shapes and blurry appearance for different subjects (Fig. 4, rows 1 and 4). SunStage assumes a single point light in the scene, and the shadows and specular components come only from this point light. This leads to poor diffuse and specular separation and incorrect shadows in generic environments like the examples shown here (Fig. 4, rows 2 and 5). In the first subject, we see orange artifacts on the forehead of the diffuse albedo in the SunStage result (row 2). In the second subject, the specular component is completely missing (row 5). The shape and textures from SunStage also have lower resolution than ours. The final row of Fig. 4 illustrates the appearance details for zoomed-in regions shown by the red and blue squares, indicating that our method produces the most accurate details.
Figure 4: Qualitative comparisons with NextFace [16] and SunStage [57] on two different subjects. The first column is the resulting render overlaid on one input image, and the remaining columns indicate the recovered appearance maps. NextFace produces inaccurate shapes and blurry appearance, where SunStage produces poor shadows and incorrect diffuse/specular separation. Our method produces the most accurate results, also indicated by the zoom region in the final row.
As a quantitative comparison, we show reconstruction errors of to NextFace, SunStage and FLARE compared to ours in Table 1. For fairness, we compute errors only on skin regions and average over all subjects in the dataset. Our method prevails in all metrics. The metrics are computed in linear RGB space and averaged over all tested subjects.
PSNR $\uparrow$ MAE $\downarrow$ SSIM [58] $\uparrow$ LPIPS [66] $\downarrow$
NextFace [16] 25.30 10.63 0.78 0.31
SunStage [57] 29.47 5.28 0.88 0.14
FLARE [5] 30.40 2.01 0.94 0.15
Ours 38.09 1.18 0.97 0.10
Table 1: Reconstruction errors computed over the skin region averaged for all the subjects.
4.3 Ablations
Raytracing vs pre-filtered environment map for diffuse.
We first compare ray tracing versus pre-filtered environment map in terms of joint reconstruction of diffuse albedo, geometry and illumination. The specular component is discussed separately in the next paragraph as it contributes mostly to the micro- instead of the macro-geometry. We created a synthetic dataset with a monocular head rotation sequence similar to the real data. The ground truth mesh and albedo are from a studio appearance capture method [52] (Fig. 5 column 1). We use a natural outdoor environment map on a cloudy day as lighting. Please refer to the supplementary document for more details on the synthetic dataset. We show reconstructed diffuse albedo and geometry of ray tracing and pre-filtered environment map in columns 2 and 3 of Fig. 5, and the error maps compared with the ground truth are shown in the last two columns. We can see that ray tracing produces a more accurate reconstruction of the diffuse albedo while the result from the pre-filtered environment map contains a lot of baked-in shading. In terms of shape reconstruction, the two perform similarly. However, the pre-filtered environment map does not handle shadows well, leading to slightly worse shape recovery in regions where self-occlusion plays an important role, e.g., the lips.
Raytracing vs pre-filtered environment map for specular.
The original render from the pre-filtered environment map has some unrealistic highlights at the grazing angle on the side of the nose when inspecting the specular component (Fig. 6, row 1 column 2). This is because the shading model does not account for light visibility and shows the contribution of the light from behind the head. Our proposed view-dependent visibility attenuation (Fig. 6, row 1 column 1) effectively masks out these regions and the final render matches the ray traced version closely. However, in our inverse rendering setting when we need to solve for shape, textures, lighting at the same time, ray tracing produces flickery images (Fig. 6 row 2) where part of the skin abruptly changes from very bright to very dark while the head is rotating. We also notice high frequency artifacts, e.g. sharp boundary between bright and dark pixels, in the ray traced render, as denoted by the red arrows. In contrast, our visibility modulated pre-filtered environment maps used for the specular component give visually smoother specular renders (Fig. 6, row 3). Note that the ray traced and pre-filtered environment map results are rendered from different sets of optimized lighting and textures in this example so the specular highlights do not match exactly.
Figure 5: Comparing ray tracing and pre-filtered environment map on a synthetic dataset rendered with a Lambertian diffuse material. The mesh errors are displayed with a scale of 0mm Refer to caption 5mm and the albedo error with a scale of -0.1 Refer to caption 0.1.
Figure 6: Comparing the specular renders from ray tracing and a pre-filtered environment map. Our proposed view-dependent specular visibility term correctly masks out the unrealistic highlights on the side of the nose, resembling the ray traced results, while produces smoother renders.
Regularize diffuse component.
We apply a weak regularization to encourage the diffuse render to be small which prevents too much specular from being baked into the diffuse component, as we show in Fig. 7. This gives us better separation of the diffuse and specular signals.
Figure 7: Diffuse and specular component without and with the regularization term $\mathcal{L}_{\text{diffuse}}$.
Optimizing vertex positions vs blendweights.
Last, we choose to optimize vertex positions directly instead of parameterizing the shape using blendweights of a 3DMM as in NextFace [16]. We show in Fig. 8 that our model achieves the least reconstruction error in the central part of the face compared to the initial mesh or the one optimized through a 3DMM in the same inverse rendering setting. Even though the silhouette of the side face is never shown in the training data, our method recovers the correct shape of the nose from shading. The ground truth reference mesh is a 3D facial scan from a multi-view face scanner [4].
Figure 8: Our geometry optimization pipeline recovers the correct nose shape from shading, achieving the better reconstruction quality compared to 3DMM fit. The mesh errors are displayed with a scale of 0mm Refer to caption 5mm
5 Limitations
We assume the head poses (and neck rotations) are provided as input to our method. Inaccuracies in the head pose estimation impair the reconstruction quality of our method substantially (see failure cases in the supplemental). Although we make no assumptions on the lighting condition, we cannot recover the appearance if part of the face is in extreme shadows in all frames. Our current face template geometry does not model eyes, however switching to a different template model would allow to reconstruct eyes better. Also, correct skin tone recovery is not guaranteed due to the ambiguity between the illumination and appearance.
6 Conclusion
We propose a new lightweight facial appearance capture method, surpassing the quality of existing lightweight approaches. Our method works truly in the wild, making no assumption on the environment illumination. Our novel shading model that combines ray tracing and a visibility modulated environment map leads to faithful recovery of the shape and appearance properties from a monocular video containing a simple head rotation.
Supplementary Material
In this supplementary document we begin by discussing the details of our capture protocol and initial tracking steps in Section 7. In Section 8 we highlight the particulars of our implementation of the baseline methods we use for comparisons, and elaborate on the synthetic dataset we used for evaluation. Finally, we add additional results and failure cases in Section 9.
7 Dataset Details
7.1 Capture Protocol
We record videos at 25 fps using a Canon EOS 1200D camera fixed on a tripod. Depending on the distance to the subject, the camera is mounted with either a 35mm or 60mm lens and we use the corresponding focal length as a known parameter when calibrating the camera. The subjects are asked to slowly rotate their head 20 to 30 degrees to the left, right, up, and down directions. We do not record large head rotations as we noticed that the estimated head poses tend to be inaccurate on side view. Fig. 9 (row 1) shows some example frames of a dataset, where the left-most and right-most frames are the most extreme head rotations in this sequence. Note that even though we do not see a side face silhouette in the data, our method can recover the correct nose shape from shading as we have shown in Fig. 8. The datasets we captured span across multiple days at different locations.
7.2 Initial Tracking
In the initial tracking stage, we estimate an initial mesh and per frame head poses. The initial mesh is parameterized using blendweights of a PCA face basis computed from the dataset of Chandran et al. [11], which includes 50 eigen faces for identity and 25 for expression. Our tracking algorithm uses a combination of a landmark [12] loss and a photometric loss, similar to Qian [50]. The only difference is that we solve for a global expression code since we assume the expression does not change in the same sequence. We apply a weight of 100 on the landmark loss and a weight of 30 on the photometric loss for all our datasets.
We further obtain an initial albedo estimate by averaging the projected texture across all the frames, and then we compute a piece-wise constant version based on a predefined face segmentation as in Rainer et al. [51]. The specular intensity map is initialized as a grayscale version of the initial diffuse albedo.
Figure 9: Example frames and initial tracking of a dataset. The first row shows in the input frames and the second row is the initial tracked geometry overlaid on the input.
8 Experiment Details
8.1 Baseline Implementation
Next we describe the steps performed to run the baseline comparisons. We use the default parameters of FLARE [5]. We noticed that the FLARE geometry is very bumpy, hence we tried setting a larger weight on the Laplacian mesh regularizer. However, this resulted in flatter face geometry and did not improve the results.
For NextFace [16, 17, 18], we obtained the best results in our experiments using only three frames that cover the whole face region. A similar behavior was observed by Azinović et al. [2] in their experiments.
The original capture protocol used in SunStage [57] has a different format compared to ours; they record only a frontal view with the person rotating 360 degrees in place. We thus adapt the preprocessing code to the one from FLARE when running SunStage. We also do not solve for the focal length of the camera and set it as the ground truth value. In our experiments, the shape does not change much from the initial DECA [22] result, in both the coarse alignment stage and the photometric optimization stage of SunStage.
NextFace and SunStage have no code for relighting in their release, so we did not compare relighting performance against these two baselines. The statistics in Table 1 are averaged over frames used in the optimization, i.e., all the frames for our method, FLARE, SunStage, and only three frames for NextFace.
8.2 Synthetic Dataset
Figure 10: Assets and example frames of the synthetic dataset.
We render a synthetic dataset with Lambertian material for the ablation study in Fig. 5. The assets, i.e. ground truth diffuse albedo, mesh and environment maps are shown in row 1 of Fig. 10, and example frames are shown in row 2. The generated head poses are similar to those from a real dataset. We perform the same initial tracking algorithm using only the landmark loss on a front-facing frame. Instead of solving per frame head poses as for a real dataset, we use the ground truth head poses for the synthetic dataset in the inverse rendering stage.
9 Additional Results
Next we present additional results for the ablation and show some challenging situations for our algorithm.
Visualization of specular visibility.
First we provide additional visualizations of the view-dependent specular visibility under different roughness in Fig. 11. This highlights the areas that are impacted by the visibility computation. When the roughness is small (mirror material), this visibility term is close to a binary mask and when the roughness is large (diffuse material), it gets closer to a view-independent ambient occlusion term. Note that the approximation error of Eq. 7 is small for a smaller roughness value and big for a larger roughness value. Applying the same approximation for the diffuse component would lead to a very large error.
Figure 11: Visualization of the estimated view-dependent specular visibility term with difference roughness values.
9.1 Failure Cases
One of the limitations of our method is that we rely on good head pose estimation from the initial tracking stage. If the initial tracking is incorrect or imprecise, see (Fig. 12 row 2), a misalignment of the render and the input image (Fig. 12 row 3 column 3) occurs, making the face appear distorted. We tried optimizing head poses in the inverse rendering stage but the results are often jittery over time and the textures are blurrier. Therefore, we decided to rely solely on the initial tracking for the head pose. Note however, that our method still produces reasonable results even when the head poses are inaccurate. In these cases, our method explains the discrepancies between the tracked mesh and the input image using texture.
Although in theory our method can work under arbitrary static lighting conditions, there are challenging cases when our method still does not produce a good enough appearance decomposition. One such example is shown in Fig. 13. In this example, the left side of the face is overexposed while the right side is much darker in all frames. While the render still matches the input image, the reconstructed diffuse albedo and specular intensity maps contain a considerable amount of baked-in lighting. Note however, that our model still manages to disentangle a major part of the lighting from the appearance, i.e. the brightness on the left and right sides of the diffuse albedo is similar. Capturing the same subject under multiple different lighting conditions can potentially improve the disentanglement [3], which we leave as future work.
Figure 12: Results from good and bad head pose estimation from the initial tracking stage. The render error maps are displayed with a scale of -0.05 Refer to caption 0.05.
Figure 13: Poor appearance decomposition in challenging lighting conditions. While the render still matches the ground truth image, the albedo map contains some minor baked-in lighting.

### v2

@@
 As such, in today’s high-end facial scanning pipelines the desired result includes a high-resolution facial surface mesh with corresponding appearance textures for properties like the diffuse albedo, specular intensity and specular roughness, which are compatible with modern skin shaders.
-Like facial geometry reconstruction, the field of skin appearance estimation is also well-studied in controlled studio environments, where accurate appearance maps can be reconstructed from large setups that obtain multi-view or multi-shot images under calibrated lighting.
+Like facial geometry reconstruction, the field of skin appearance estimation is also well-studied in controlled studio environments, where accurate appearance maps can be reconstructed from large setups that obtain multi-view or multi-shot images under calibrated lighting [55].
 Following the geometry trend, current research aims to allow facial appearance capture in less constrained settings, for example outdoors using the sun as a single point light [62].
@@
 Different from previous methods, we do not make any assumptions on the lighting condition (e.g.
-we do not require a sun in the sky), and we explicitly take visibility into account, effectively removing baked-in shading by correctly modeling self-occlusion in our appearance solver.
-We do so by proposing a novel shading model that combines ray tracing and a visibility modulated pre-filtered environment map.
+we do not require a sun in the sky), and as our main contribution we explicitly take visibility into account, effectively removing baked-in shading by correctly modeling self-occlusion in our appearance solver.
+The result is a detailed geometry mesh with textures for diffuse albedo, specular intensity and roughness.
 As we will show, our approach leads to more faithful recovery of the appearance properties than existing techniques in the wild.
-The result is a detailed geometry mesh with textures for diffuse albedo, specular intensity and roughness, which approach the quality achievable from studio-based setups but are captured more flexibly in the wild.
 As a particular application, our approach allows fast capture of actors on a film set, with resulting assets that can be used directly in traditional VFX pipelines.
@@
   • A new state-of-the-art method for in-the-wild facial appearance capture that makes no assumption on the scene lighting condition.
-  • A novel shading model that explicitly handles self-occlusion by combining ray tracing with pre-filtered environment maps, achieving high-quality appearance reconstruction from monocular input.
+  • A novel shading model that explicitly handles visibility and self-occlusion for inverse rendering pipelines, achieving high-quality appearance reconstruction from monocular input.
 2 Related Work
@@
 PhySG [70] utilizes spherical Gaussians to approximately and efficiently evaluate the rendering equation in closed form.
-Munkberg et al.
-[46] propose an efficient end-to-end framework for joint learning of topology, triangle meshes and materials, achieving much faster training and inference compared to previous NeRF-based factorization methods [67, 8, 55].
+Munkberg et al[48] propose an efficient end-to-end framework for joint learning of topology, triangle meshes and materials, achieving much faster training and inference compared to previous NeRF-based factorization methods [72, 7, 59].
 They also introduce a differentiable formulation of the split-sum approximation of environment lighting to efficiently recover all-frequency lighting.
@@
 Similar setups were used for facial appearance decomposition with neural and gaussian primitives [56, 69, 67, 66, 57].
-Lighter alternative setups were also explored by Lattas et al.
-[37] and Choi et al.
-[14] but they still require the subject to be seating in a dedicated space.
+Lighter alternative setups were also explored by Lattas et al[39] and Choi et al[13] but they still require the subject to be seating in a dedicated space.
 Recently, the research community has investigated more lightweight setups which are easily accessible to everyone.
@@
 CoRA [28] reconstructs relightable 3D face assets from a single co-located smartphone flashlight sequence captured in a dim room.
-Similarly, Azinović et al.
-[2] additionally attach polarization foils and capture two such sequences with perpendicular polarization orientation to separate skin surface and subsurface reflectance.
+Similarly, Azinović et al[1] additionally attach polarization foils and capture two such sequences with perpendicular polarization orientation to separate skin surface and subsurface reflectance.
 Using a co-located light and camera setup, these methods assume the position of the dominant light source is known and no shadowing term needs to be modeled.
@@
 All these methods assume the specular reflection (and the shadowing) of the face comes from a single dominant point or directional light source in the capture environment, and the ambient light contributes only to a low-frequency diffuse term.
-NextFace [16, 17, 18] exploits a differentiable ray tracer to estimate various scene attributes given unconstrained monocular images.
-It relies heavily on statistical shape and appearance priors [7, 24, 54] which limits its expressiveness.
+Another line of work [15, 16, 17, 38, 27, 51, 18] aims to reconstruct shape and reflectance properties from a single portrait image, often relying heavily on statistical shape and appearance priors [6, 24, 58] which limits its expressiveness.
 Our work instead focuses on accurate personalized reconstruction.
-Rainer et al.
-[51] use tiny shading networks to disentangle shading from explicit reflectance maps but assume a known high-quality geometry and smooth lighting.
+Rainer et al[54] use tiny shading networks to disentangle shading from explicit reflectance maps but assume a known high-quality geometry and smooth lighting.
 NeuFace [73] represents the face with a neural SDF and proposes to learn appearance factorization under unknown low frequency light with a novel neural BRDF basis.
@@
 Closest to ours, FLARE [4] builds relightable head avatars from monocular videos.
-It uses pre-filtered environment maps for relighting and a neural approximation of it during training.
+It adopts the split-sum approximation [32, 48] for relighting and a neural version of it during training.
 This rendering model ignores the self-occlusion and bakes part of the shading into the albedos.
-In contrast, we explicitly handle light visibility, using ray tracing for diffuse reflectance and visibility modulated pre-filtered environment map for the specular reflectance, leading to higher quality shape and appearance reconstruction.
+In contrast, we propose a modified formulation of the split-sum approximation which explicitly handles light visibility and combine it with ray tracing, leading to higher quality shape and appearance reconstruction.
 3 Monocular Appearance Capture
@@
 The output of our inverse rendering system will be a 3D asset of the subject containing a high-quality mesh, and diffuse albedo, specular intensity and roughness as 2D texture maps.
-In the following, we first describe our geometry optimization formulation in Section 3.1, shading model in Section 3.2, and then implementation details in Section 3.3.
+In the following, we first describe our geometry optimization formulation in Section 3.1, our novel occlusion-aware shading model in Section 3.2, and then optimization details in Section 3.3.
 3.1 Geometry Optimization
@@
 Our method, however, optimizes geometry and textures at the same time.
-To do so, we adopt a preconditioning framework similar to Nicolet et al.
-[48], which biases gradient steps towards smooth solutions.
+To do so, we adopt a preconditioning framework similar to Nicolet et al[50], which biases gradient steps towards smooth solutions.
 The vertex positions $\mathbf{v}$ in each iteration are updated by
@@
 This way, we can apply a large learning rate for the geometry optimization while keeping the mesh smooth and self-intersection free.
-3.2 Shading Model
-Following the rendering equation [30], we compute the outgoing radiance $L(\omega_{o})$ in direction $\omega_{o}$ by:
-[…]
-We decompose the BRDF $f(\mathbf{x},\omega_{i},\omega_{o})$ into a diffuse term $f_{d}$ and a specular term $f_{s}$. We use the simple Lambertian model for the diffuse term:
+3.2 Occlusion-Aware Shading Model
+Following the rendering equation [31], we compute the outgoing radiance $L(\omega_{o})$ at location $\mathbf{x}$ from direction $\omega_{o}$ by:
+$L(\omega_{o})=\int_{\Omega}f(\mathbf{x},\omega_{i},\omega_{o})L_{i}(\omega_{i})(\omega_{i}\cdot\mathbf{n})d\omega_{i}.$ (2)
+We decompose the BRDF $f(\mathbf{x},\omega_{i},\omega_{o})$ into the sum of a diffuse term $f_{d}$ and a specular term $f_{s}$. We use the simple Lambertian model for the diffuse term:
 $f_{d}(\mathbf{x})=\frac{\rho(\mathbf{x})}{\pi},$ (3)
-where $\rho$ is the diffuse albedo. We use a specular BRDF similar to Kelemen and Szirmay-Kalos [34], which has been shown to be well suited for rendering human skin [47, Chapter 14]:
+where $\rho$ is the diffuse albedo. We use a specular BRDF similar to Kelemen and Szirmay-Kalos [35], which has been shown to be well suited for rendering human skin [49, Chapter 14]:
 $f_{s}(\mathbf{x},\omega_{i},\omega_{o})=\frac{DGF}{4(\omega_{i}\cdot\mathbf{n})(\omega_{o}\cdot\mathbf{n})},$ (4)
 where $D$, $G$, and $F$ are functions representing the Beckmann normal distribution, geometric attenuation and Fresnel terms, respectively.
-Accounting for light visibility.
+Accounting for self-occlusion.
 We consider direct illumination where $L_{i}(\omega_{i})$ comes only from the light sources.
 Spherical Harmonics [44, 12] or Spherical Gaussians [68] are popular representations used by prior methods [70, 73] but they can only model low- to medium-frequency lighting.
-Other work [5, 46] uses pre-filtered environment maps [32, 43], which allows to capture all-frequency lighting but still ignores self-shadowing.
-We aim to use a more realistic rendering model that accounts for light visibility to achieve higher photorealism.
-For the diffuse term, we apply explicit ray tracing to compute the outgoing radiance and use
-[…]
-where $V(\mathbf{x},\omega_{i})$ is the visibility at surface point $\mathbf{x}$ from incoming light direction $\omega_{i}$.
-We perform Monte Carlo integration with multiple importance sampling similar to Hasselgren et al.
-[28] skipping the denoising step as we did not observe any noise issue for the diffuse term in our experiments.
-We empirically found that ray tracing for the specular term tends to produce flickery renders (Figure 6) due to our challenging scenario where only one camera view is available.
-Therefore, we propose to use a pre-filtered environment map but modulate the specular render with a view-dependent visibility term.
-We apply a split-sum approximation similar to Karis [31] and use
-$L_{\text{specular}}(\omega_{o})\approx$ (6)
-[…]
-The first term is the material integral that depends on $\omega_{i}\cdot\mathbf{n}$ and the roughness.
-The second integral is the pre-filtered environment map.
-Both integrals can be pre-computed and stored as 2D textures.
-At rendering time, the rendering integral can be evaluated using only two texture lookups.
-When light visibility is not accounted for, the second integral only depends on the reflected view direction $\omega_{r}$ and the roughness $r$.
-When $r\ll 1$, the second integral is 0 unless $\omega_{i}$ is close to $\omega_{r}$.
+We therefore follow other work, e.g.
+FLARE [4], and use a differentiable split-sum approximation [33, 45, 48], which allows to capture all-frequency lighting.
+Unfortunately this approach ignores self-shadowing, and so we propose a novel visibility-modulated split-sum approximation to account for self-shadowing, which we describe in the following.
+We start by introducing the split-sum formulation from Karis [32] which approximates  Eq. 2 as
+$L(\omega_{o})\approx$ (5)
+$\int_{\Omega}f(\mathbf{x},\omega_{i},\omega_{o})(\omega_{i}\cdot\mathbf{n})d\omega_{i}\int_{\Omega}L_{i}(\omega_{i})D(\mathbf{h})(\omega_{i}\cdot\mathbf{n})d\omega_{i}.$
+The first term is the integral of the BRDF under a solid white environment map, i.e.
+$L_{i}(\omega_{i})=1,\forall\omega_{i}$.
+It only depends on $\cos\theta=\omega_{i}\cdot\mathbf{n}$ and the roughness $r$.
+Therefore, it can be precomputed and stored as a 2D look-up texture.
+The second integral is a pre-filtered environment map, where $D$ is the normal distribution function of the BRDF and $\mathbf{h}=\frac{\omega_{i}+\omega_{o}}{\|\omega_{i}+\omega_{o}\|_{2}}$ is the half vector.
+This term can also be precomputed as a mipmap by convolving the environment map with $D$ at different roughness values.
+At rendering time, the outgoing radiance can then be efficient computed using only two texture lookups.
+Note that when evaluating  Eq.
+5, the contribution of the prefiltered environment map is only dependent on the reflected view direction $\omega_{r}$ and the roughness $r$.
+This is not physically accurate as the contribution should be reduced if the light is (partially) blocked from the upper hemisphere at a specific point.
+To account for self-occlusion, we must modulate $L_{i}(\omega_{i})$ differently at different locations based on the light visibility $V(\mathbf{x},\omega_{i})$, i.e., making the second integral in  Eq.
+5 $\int_{\Omega}L_{i}(\omega_{i})V(\mathbf{x},\omega_{i})D(\mathbf{h})(\omega_{i}\cdot\mathbf{n})d\omega_{i}$.
+This new integral cannot be precomputed anymore since it’s different for different $\mathbf{x}$.
+However, we notice that when $r\ll 1$, the new integral is 0 unless $\omega_{i}$ is close to $\omega_{r}$.
 Therefore, we can approximate it as
 Figure 1: A selection of facial appearance reconstruction and decomposition results for different subjects in different environments, both indoors and outdoors, with sunny and cloudy sky.
-[…]
-[…]
-Here, $\tilde{V}(\mathbf{x},\omega_{r})$ is the view-dependent specular visibility.
+$\int_{\Omega}L_{i}(\omega_{i})V(\mathbf{x},\omega_{i})D(\mathbf{h})(\omega_{i}\cdot\mathbf{n})d\omega_{i}$ (6)
+$\approx$ $\tilde{V}(\mathbf{x},\omega_{r})\int_{\Omega}L_{i}(\omega_{i})D(\mathbf{h})(\omega_{i}\cdot\mathbf{n})d\omega_{i}.$
+Here, $\tilde{V}(\mathbf{x},\omega_{r})$ is the view-dependent visibility.
 The simplest choice of $\tilde{V}(\mathbf{x},\omega_{r})$ is to only evaluate the light visibility at $\omega_{r}$, i.e., $\tilde{V}(\mathbf{x},\omega_{r})\coloneqq V(\mathbf{x},\omega_{r})$.
-Intuitively, this means the specular component is 0 if the reflected view direction is occluded.
+Intuitively, this means the outgoing radiance is 0 if the reflected view direction is occluded.
 Note that this approximation is exact for perfect specular reflection (i.e.
 mirrors).
-To add some softness in the visibility attenuation term (instead of considering it as a binary function), we approximate it by Monte Carlo integration,
-$\tilde{V}(\mathbf{x},\omega_{r})\coloneqq\frac{1}{K}\sum_{k=1}^{K}\frac{V(\mathbf{x},\omega_{k})}{p(\omega_{k})},$ (8)
-where the samples are drawn following the normal distribution of the specular lobe.
-This bears some resemblance to the appearance models of Saito et al.
-[53].
-However, they use Spherical Gaussians for the specular lobe and parameterize the view-dependent specular visibility term using a neural network as they work with a large amount of studio data where the lighting is controlled and calibrated.
-3.3 Implementation Details
-Our training loss consists of an $L_{1}$ data term $\mathcal{L}_{\text{img}}$ and some regularization terms.
+To add some softness in the visibility term (instead of considering it as a binary function), we approximate it by Monte Carlo integration,
+$\tilde{V}(\mathbf{x},\omega_{r})\coloneqq\frac{1}{K}\sum_{k=1}^{K}\frac{V(\mathbf{x},\omega_{k})}{D(\mathbf{n},\omega_{k},\omega_{r},r)},$ (7)
+where the samples are drawn following the normal distribution of the BRDF.
+This bears some resemblance to the appearance models of Saito et al[56].
+However, they parameterize the view-dependent specular visibility term using a neural network as they work with a large amount of studio data where the lighting is controlled and calibrated.
+Note however,  Eq.
+6 introduces large errors for rough surfaces (i.e., when $r\ll 1$ does not hold).
+We therefore use it only for the specular component, and implemented ray tracing with multiple importance sampling [60] for the diffuse component using the OptiX [52] engine.
+3.3 Optimization Details
+Our reconstruction loss consists of an $L_{1}$ data term $\mathcal{L}_{\text{img}}$ and some regularization terms.
 We employ an $L_{1}$ mask loss $\mathcal{L}_{\text{mask}}$ between the mask obtained from MODNet [34] and the predicted binary mask.
-Although we use a parameterization similar to Nicolet et al.
-[48], we find it helpful to still employ a Laplacian regularizer to stabilize the geometry optimization such that we do not need to set different $\lambda_{\text{geo}}$ for different datasets.
+Although we use a parameterization similar to Nicolet et al[50], we find it helpful to still employ a Laplacian regularizer to stabilize the geometry optimization such that we do not need to set different $\lambda_{\text{geo}}$ for different datasets.
 This regularizer encourages the Laplacian of the optimized mesh to stay close to the Laplacian of the initial 3DMM fit
@@
 $\mathcal{L}\coloneqq$ $\mathcal{L}_{\text{img}}+\lambda_{\text{mask}}\mathcal{L}_{\text{mask}}+\lambda_{\text{Lap}}\mathcal{L}_{\text{Lap}}+$ (10)
-$\lambda_{\text{light}}\mathcal{L}_{\text{light}}+\lambda_{\text{rough}}\mathcal{L}_{\text{rough}}+\lambda_{\text{diffuse}}\mathcal{L}_{\text{diffuse}},$
-where the weights are defined as
-$\lambda_{\text{mask}}=0.1,\lambda_{\text{Lap}}=10,\lambda_{\text{light}}=0.1,$ (12)
-$\lambda_{\text{rough}}=0.1,\lambda_{\text{diffuse}}=0.01.$
-We capture 500 to 800 frames for each subject.
-Please see the supplemental PDF for details of our capture protocol.
-We then uniformly sample around 250 frames to use in the inverse rendering.
-We did not observe a performance gain or drop when using all of the frames.
-The images are cropped to 1K resolution.
-We also solve the texture maps at 1K resolution.
-The environment map is a cubemap with resolution $6\times 256\times 256$ at the largest mip-level, and resolution $6\times 8\times 8$ at the smallest mip-level.
-For each pixel, we draw $256$ light samples and $256$ BRDF (cosine) samples for the diffuse render, and $64$ samples to estimate the view-dependent specular visibility.
-We employ the Adam optimizer [19] with a learning of $0.1$ for vertex positions and the environment map, and $0.001$ for the textures.
-We use the differentiable rasterizer from Laine et al.
-[36] to obtain the primary visibility and the OptiX [49] engine for ray tracing.
-Each subject is trained for $6000$ iterations which takes around 2 hours on a Nvidia RTX 3090 GPU.
+$\lambda_{\text{light}}\mathcal{L}_{\text{light}}+\lambda_{\text{rough}}\mathcal{L}_{\text{rough}}+\lambda_{\text{diffuse}}\mathcal{L}_{\text{diffuse}}.$
 4 Experiments
@@
 1 for the original environment).
+Please refer to the supplemental video for more results.
 Figure 2: Relighting results of multiple frames on two different subjects in two different environments. The environment is shown top left of each row. See Fig. 1 for the original environment.
@@
 4, rows 1 and 4).
+It produces very similar shapes for the two subjects shown in Fig.
+4, while our geometry preserves the identity and likeness of the subjects.
 SunStage assumes a single point light in the scene, and the shadows and specular components come only from this point light.
@@
 FLARE [4] 30.40 2.01 0.94 0.15
+Ours (w/o vis) 34.55 1.79 0.96 0.10
 Ours 38.09 1.18 0.97 0.10
 Table 1: Reconstruction errors computed over the skin region averaged for all the subjects.
-4.3 Ablations
-Raytracing vs pre-filtered environment map for diffuse.
-We first compare ray tracing versus pre-filtered environment map in terms of joint reconstruction of diffuse albedo, geometry and illumination.
-The specular component is discussed separately in the next paragraph as it contributes mostly to the micro- instead of the macro-geometry.
-We created a synthetic dataset with a monocular head rotation sequence similar to the real data.
+4.3 Evaluations
+Effects of light visibility.
+We first evaluate the effects of accounting for self-occlusion by comparing the relit renders of our method with the original split-sum approximation (denoted as Ours (w/o vis)) in Fig.
+5.
+We can see that when self-occlusion is not accounted for, shadows under the capture lighting got baked into the albedo (column 1 with zoomed-in patches focusing around the nose), leading to wrong shadows when relit (row 1, column 2, 3).
+The specular component also exhibits unrealistic sharp highlight (row 1, column 2, 4).
+Moreover, when the major light source is behind the subject, the baseline renders show artifacts in the form of strong specular highlights on the side of the nose and under the chin (row 3, column 2, 4) since self-occlusion is not properly handled.
+In contrast, our proposed model correctly removes baked-in shading and produces more realistic relit renders (row 2, 4).
+We also quantitatively evaluated the performance gain of accounting for self-occlusion in terms of reconstruction errors in Table 1.
+Figure 5: Relighting evaluation of our method with and without accounting for self-occlusion. We also visualize the corresponding diffuse albedo, diffuse render and specular render.
+Evaluation on synthetic data.
+To better evaluate the quality of the reconstructed mesh and textures against ground truth, we created a synthetic dataset with a monocular head rotation sequence similar to the real data.
 The ground truth mesh and albedo are from a studio appearance capture method [55] (Fig.
@@
 Please refer to the supplementary document for more details on the synthetic dataset.
-We show reconstructed diffuse albedo and geometry of ray tracing and pre-filtered environment map in columns 2 and 3 of Fig.
-5, and the error maps compared with the ground truth are shown in the last two columns.
-We can see that ray tracing produces a more accurate reconstruction of the diffuse albedo while the result from the pre-filtered environment map contains a lot of baked-in shading.
+We show reconstructed diffuse albedo and geometry of our method with and without light visbility accounted for in columns 2 and 3 of Fig.
+6, and the error maps compared with the ground truth are shown in the last two columns.
+We can see that our method produces a more accurate reconstruction of the diffuse albedo while the baseline result contains a lot of baked-in shading.
 In terms of shape reconstruction, the two perform similarly.
-However, the pre-filtered environment map does not handle shadows well, leading to slightly worse shape recovery in regions where self-occlusion plays an important role, e.g., the lips.
-Raytracing vs pre-filtered environment map for specular.
-The original render from the pre-filtered environment map has some unrealistic highlights at the grazing angle on the side of the nose when inspecting the specular component (Fig.
-6, row 1 column 2).
-This is because the shading model does not account for light visibility and shows the contribution of the light from behind the head.
-Our proposed view-dependent visibility attenuation (Fig.
-6, row 1 column 1) effectively masks out these regions and the final render matches the ray traced version closely.
-However, in our inverse rendering setting when we need to solve for shape, textures, lighting at the same time, ray tracing produces flickery images (Fig.
+However, the baseline method does not handle shadows well, leading to slightly worse shape recovery in regions where self-occlusion plays an important role, e.g., the lips.
+Ours vs raytracing for specular.
+While a ray-tracer handles visibility inherently, we find that our modified split-sum approximation behaves more stably in our inverse rendering setting when we need to solve for shape, textures, lighting at the same time from merely monocular input.
+Ray tracing produces flickery specular images (Fig.
 7 row 2) where part of the skin abruptly changes from very bright to very dark while the head is rotating.
@@
 sharp boundary between bright and dark pixels, in the ray traced render, as denoted by the red arrows.
-In contrast, our visibility modulated pre-filtered environment maps used for the specular component give visually smoother specular renders (Fig.
-6, row 3).
-Note that the ray traced and pre-filtered environment map results are rendered from different sets of optimized lighting and textures in this example so the specular highlights do not match exactly.
-Figure 5: Comparing ray tracing and pre-filtered environment map on a synthetic dataset rendered with a Lambertian diffuse material.
+In contrast, our model gives visually smoother specular renders (Fig.
+7, row 3).
+Please refer to the supplemental videos for an animated visualization.
+Figure 6: Comparing our method with and without accounting for light visibility on a synthetic dataset.
 The mesh errors are displayed with a scale of 0mm Refer to caption 5mm and the albedo error with a scale of -0.1 Refer to caption 0.1.
-Figure 6: Comparing the specular renders from ray tracing and a pre-filtered environment map.
-Our proposed view-dependent specular visibility term correctly masks out the unrealistic highlights on the side of the nose, resembling the ray traced results, while produces smoother renders.
-Regularize diffuse component.
-We apply a weak regularization to encourage the diffuse render to be small which prevents too much specular from being baked into the diffuse component, as we show in Fig.
-7.
-This gives us better separation of the diffuse and specular signals.
-Figure 7: Diffuse and specular component without and with the regularization term $\mathcal{L}_{\text{diffuse}}$.
+Figure 7: Comparing the specular renders from our method and a ray-tracer across frames in the same head rotation sequence.
 Optimizing vertex positions vs blendweights.
-Last, we choose to optimize vertex positions directly instead of parameterizing the shape using blendweights of a 3DMM as in NextFace [16].
+We choose to optimize vertex positions directly instead of parameterizing the shape using blendweights of a 3DMM as in NextFace [15].
 We show in Fig.
@@
 Figure 8: Our geometry optimization pipeline recovers the correct nose shape from shading, achieving the better reconstruction quality compared to 3DMM fit.
-The mesh errors are displayed with a scale of 0mm Refer to caption 5mm
+The mesh errors are displayed with a scale of 0mm Refer to caption 5mm.
+Regularize diffuse component.
+Last, we apply a weak regularization to encourage the diffuse render to be small which prevents too much specular from being baked into the diffuse component, as we show in Fig.
+9.
+This gives us better separation of the diffuse and specular signals.
+Figure 9: Diffuse and specular component without and with the regularization term $\mathcal{L}_{\text{diffuse}}$.
 5 Limitations
@@
 We propose a new lightweight facial appearance capture method, surpassing the quality of existing lightweight approaches.
-Our method works truly in the wild, making no assumption on the environment illumination.
-Our novel shading model that combines ray tracing and a visibility modulated environment map leads to faithful recovery of the shape and appearance properties from a monocular video containing a simple head rotation.
+It works truly in the wild, making no assumption on the environment illumination.
+Our novel shading model explicitly accounts for self-occlusion, leading to faithful recovery of the shape and appearance properties from a monocular video containing a simple head rotation.
 Supplementary Material
 In this supplementary document we begin by discussing the details of our capture protocol and initial tracking steps in Section 7.
-In Section 8 we highlight the particulars of our implementation of the baseline methods we use for comparisons, and elaborate on the synthetic dataset we used for evaluation.
+We provide more implementation details in Section 8.
+In Section 9 we highlight the particulars of our implementation of the related methods we use for comparisons, and elaborate on the synthetic dataset we used for evaluation.
 Finally, we add additional results and failure cases in Section 10.
@@
 [10], which includes 50 eigen faces for identity and 25 for expression.
-Our tracking algorithm uses a combination of a landmark [12] loss and a photometric loss, similar to Qian [50].
+Our tracking algorithm uses the combination of a landmark [11] loss and a photometric loss, similar to Qian [53].
 The only difference is that we solve for a global expression code since we assume the expression does not change in the same sequence.
@@
 Figure 10: Example frames and initial tracking of a dataset. The first row shows in the input frames and the second row is the initial tracked geometry overlaid on the input.
+8 Implementation Details
+We capture 500 to 800 frames for each subject.
+We then uniformly sample around 250 frames to use in the inverse rendering.
+We did not observe a performance gain or drop when using all of the frames.
+The images are cropped to 1K resolution.
+We also solve the texture maps at 1K resolution.
+The environment map is a cubemap with resolution $6\times 256\times 256$ at the largest mip-level, and resolution $6\times 8\times 8$ at the smallest mip-level.
+For each pixel, we draw $256$ light samples and $256$ BRDF (cosine) samples for the diffuse render, and $64$ samples to estimate the view-dependent specular visibility.
+We employ the Adam optimizer [19] with a learning of $0.1$ for vertex positions and the environment map, and $0.001$ for the textures.
+We use the differentiable rasterizer from Laine et al.
+[37] to obtain the primary visibility and the OptiX [52] engine for ray tracing.
+Each subject is trained for $6000$ iterations which takes around 2 hours on a Nvidia RTX 3090 GPU.
+The weights for Eq.
+10 are defined as
+$0.1,\displaystyle\lambda_{\text{mask}}=1,\lambda_{\text{Lap}}=0,\lambda_{\text{light}}=$1, (11)
+[…]
 9 Experiment Details
-8.1 Baseline Implementation
-Next we describe the steps performed to run the baseline comparisons.
+9.1 Implementation of the Related Methods
+Next we describe the steps performed to run the comparisons.
 We use the default parameters of FLARE [4].
@@
 We render a synthetic dataset with Lambertian material for the ablation study in Fig.
-5.
+6.
 The assets, i.e.
 ground truth diffuse albedo, mesh and environment maps are shown in row 1 of Fig.
-10, and example frames are shown in row 2.
+11, and example frames are shown in row 2.
 The generated head poses are similar to those from a real dataset.
@@
 Next we present additional results for the ablation and show some challenging situations for our algorithm.
-Visualization of specular visibility.
-First we provide additional visualizations of the view-dependent specular visibility under different roughness in Fig.
-11.
+Visualization of the Visibility.
+First we provide additional visualizations of the view-dependent visibility under different roughness in Fig.
+12.
 This highlights the areas that are impacted by the visibility computation.
@@
 6 is small for a smaller roughness value and big for a larger roughness value.
-Applying the same approximation for the diffuse component would lead to a very large error.
-Figure 11: Visualization of the estimated view-dependent specular visibility term with difference roughness values.
-9.1 Failure Cases
+Applying the same approximation for the diffuse component would lead to a large error.
+We also show an example in Fig.
+13 of how our visibility-modulated split-sum approximation can be used in other rendering tasks as a practical way to add self-shadowing to glossy objects.
+Figure 12: Visualization of the estimated view-dependent specular visibility term with different roughness values.
+Figure 13: Rendering of a glossy Happy Buddha.
+Failure Cases.
 One of the limitations of our method is that we rely on good head pose estimation from the initial tracking stage.
@@
 One such example is shown in Fig.
-13.
+15.
 In this example, the left side of the face is overexposed while the right side is much darker in all frames.
@@
 Figure 15: Poor appearance decomposition in challenging lighting conditions. While the render still matches the ground truth image, the albedo map contains some minor baked-in lighting.
+Effects of $\lambda_{geo}$.
+We use $\lambda_{geo}=19$ based on [50], but this smoothing term can be tuned for different subjects to get better geometry.
+An ablation for different values of $\lambda_{geo}$ is shown in Fig.
+16.
+Note that a large $\lambda_{geo}$ leads to loss of details, but a small $\lambda_{geo}$ can lead to self-intersections.
+Figure 16: Effects of $\lambda_{geo}$. The shaded meshes are shown on the top row with the corresponding normals visualized on the bottom.
+Novel Views in the Capture Environment.
+We show novel view renders in the capture environment in Fig.
+17.
+We primarily focused on the facial regions for this project.
+Artifacts around the boundaries can be improved with some engineering efforts, such as better masking and better initial alignment.
+Reconstruction of the hair and shoulders is also an interesting area for future work, which can improve the overall quality.
+Figure 17: Novel views rendered in the capture environment (top row) along with the corresponding meshes (bottom row).
+Comparison with CoRA.
+CoRA [28] requires a more constrained capture setup with a co-located light and camera in a dark room.
+We ran our method and CoRA on a subject using the CoRA capture protocol, see Fig.
+18, demonstrating that our method achieves equally good reconstruction given the same data.
+Additionally, our method also works in more generic environments.
+Figure 18: Comparison with CoRA on their data protocol.